Problem: This vehicle came into the shop with the MIL illuminated. The tech retrieved three DTCs: P0171 – System too lean; P0170 – Fuel trim; and P2195 – HO2S Signal stuck lean B1S1. The Long Term Fuel Trim (LTFT) was reading +49% and the Rear O2 sensor was reading rich at all times.
Details: First, the tech checked the fuel trims at higher RPM. The LTFT was still reading about +49%. The rear O2 was still reading rich (900mv). Next, he inspected the exhaust system for leaks – none found. The tech then unplugged a vacuum hose to check for changes in the LTFT readings and rear O2 voltage. The fuel trim remained high but rear O2 voltage dropped to near 100 mv.
The LTFT reading showed an extremely lean condition while the Rear O2 sensor was showing rich condition all time. Since the exhaust cannot be rich and lean at the same time, one of the O2 sensors had to be defective. Only the rear O2 sensor reacted to the vacuum leak; the front did not.
Confirmed Repair: The tech replaced the front O2 sensor, cleared the codes and retested. LTFT now read 0% to +5% and the all Emission Readiness Monitors ran to “Complete.”